# On the Expressive Power of Modern Hopfield Networks

### v1

@@ 4.1 Computing Hopfield Attention Matrix
    The term $\beta\cdot s_{i,j}$ is computed with depth $d_{\mathrm{std}}$ using Part 1 of Lemma B.5.
@@
    Using Lemma B.6, the exponential function $\exp(\beta\cdot s_{i,j})$ is computable with depth $d_{\mathrm{exp}}$.
Summing the circuit depths for all steps, the total depth required for $A_{i,j}$ is:
$d_{\mathrm{total}}=4d_{\mathrm{std}}+3d_{\oplus}+d_{\mathrm{exp}}.$
Since all entries of $A$ can be simulated in parallel, the total depth $4d_{\mathrm{std}}+3d_{\oplus}+d_{\mathrm{exp}}$ and size $\operatorname{poly}(n)$, completing the proof. ∎
@@ 4.2 Computing Single Hopfield Layer
The computation involves multiplying matrices $D^{-1},A,Y,\widetilde{W}_{V}$. First, we compute $D$ and $A$:
  1. 1.
    $D=\mathrm{diag}(\beta\cdot A\mathbf{1}_{n})$ is computed with depth $d_{\oplus}+d_{\mathrm{std}}$, as shown in Part 1 and Part 3 of Lemma B.5.
  2. 2.
    From Lemma 4.2, $A$ is computed with depth $4d_{\mathrm{std}}+3d_{\oplus}+d_{\mathrm{exp}}$.
  3. 3.
    From Lemma 4.2, $A$ is computed with depth $4d_{\mathrm{std}}+3d_{\oplus}+d_{\mathrm{exp}}$.
  4. 4.
    Finally, the division $D^{-1}\cdot(AY\widetilde{W}_{V})$ is computed in parallel with depth $d_{\mathrm{std}}$.
Combining these depths:
$d_{\mathrm{total}}=8d_{\mathrm{std}}+6d_{\oplus}+d_{\mathrm{exp}}.$
The total size of the circuit remains $\operatorname{poly}(n)$, completing the proof. ∎
@@ 4.3 Computing Common Components Layers
For each $i\in[n]$, Lemma 4.1 guarantees that the computation of $W_{1}\cdot X_{i,*}$ requires a $\mathsf{DLOGTIME}$-uniform circuit which has depth $d_{\mathrm{std}}+d_{\oplus}$ and size $\operatorname{poly}(n)$. Applying Part 1 of Lemma B.5, we can compute $W_{1}\cdot X_{i,*}+b_{1}$ with an additional $\mathsf{DLOGTIME}$-uniform circuit which has depth $d_{\mathrm{std}}$ and size $\operatorname{poly}(n)$. The ReLU activation, $\mathsf{ReLU}(W_{1}\cdot X_{i,*}+b_{1})$, also requires depth $d_{\mathrm{std}}$ and size $\operatorname{poly}(n)$ as per Part 1 of Lemma B.5.
@@ 5 Complexity of Kernelized Hopfield Networks
In this section, we extend the complexity results of modern Hopfield networks to kernelized Hopfield networks. The formal definition of kernelized Hopfield networks is provided in Appendix 3.3.
Section 5.1 analyzes the computation of the kernelized Hopfield attention matrix. In Section 5.2, we examine the computation of a single kernelized Hopfield layer. Section 5.3 details the computation of the entire kernelized Hopfield network. Finally, in Section 5.4, we present the main results on the circuit complexity bounds for kernelized Hopfield networks. In appendix E, we provide the complete proofs for this section.
@@ 5.1 Computing Kernelized Hopfield Attention Matrix
See Appendix E.1 for the complete proof. ∎
@@ 5.2 Computing Single Kernelized Hopfield Layer
See Appendix E.2 for the complete proof. ∎
@@ 5.3 Computing Kernelized Hopfield Networks
See Appendix E.3 for the complete proof. ∎
@@ 5.4 Main Result: Circuit Complexity Bound of Kernelized Hopfield Networks
See Appendix E.4 for the complete proof. ∎
@@ 6 Hardness
In this section, we present two computational problems along with their corresponding hardness results. In In Section 6.1, we outline the corresponding hardness results. Appendix D, we introduce the undirected graph connectivity problem and the tree isomorphism problem.
6.1 Hardness Results
Theorem 6.1.
Proof.
This result is obtained by integrating Theorem 4.6 (the circuit complexity bound of modern Hopfield networks), Lemma D.4 (showing that undirected graph connectivity is $\mathsf{NC}^{1}$-complete), and Fact D.3. ∎
Theorem 6.2.
Proof.
This result derives from Theorem 4.6 and Lemma D.8 (showing that tree isomorphism is $\mathsf{NC}^{1}$-complete). ∎
Theorem 6.3.
Proof.
This result is a direct consequence of Theorem 5.4 and Lemma D.4. ∎
Theorem 6.4.
Proof.
This result is a direct consequence of Theorem 5.4 and Lemma D.8. ∎
7 Conclusion
In this study, we conduct a comprehensive theoretical investigation of modern Hopfield networks and kernelized Hopfield networks, establishing key limitations on their computational power. Our analysis focuses on the circuit complexity of individual architectural components, showing that these networks can be simulated using uniform $\mathsf{TC}^{0}$ circuits. Crucially, we prove that unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$, both modern and kernelized Hopfield networks, when implemented with polynomial precision, a constant number of layers, and hidden dimensions satisfying $d\leq O(n)$, are unable to solve problems such as undirected graph connectivity and tree isomorphism. These findings highlight the inherent expressivity constraints of Hopfield models, even in light of their demonstrated empirical success across various tasks.
However, our analysis has limitations. It primarily addresses the forward computation of these networks, assuming constant-depth nonlinear activation functions, and does not consider the training process or the implications of more complex activation mechanisms. Expanding this framework to encompass other Hopfield network variants and exploring whether similar computational bounds apply to advanced architectures would be a valuable direction for future research. Additionally, our results point to an intriguing gap between the theoretical constraints and the practical effectiveness of these models. Investigating how Hopfield networks achieve strong empirical performance despite these limitations could provide deeper insights into model scaling and architecture design, paving the way for more theoretically robust approaches.
Appendix
Roadmap.
In Section A, we give an extended discussion of related works. In Section B, we introduce essential definitions for circuit complexity and float point numbers used in this paper. In Section C, we present the additional theoretical results of the modern Hopfield model with a chain of thought (CoT). In Section D, we present the formulation of graph connectivity and tree isomorphism problems. In Section E, we present the formal proofs for kernelized Hopfield networks.
Appendix A More Related Works
Modern Hopfield Networks and Applications in Deep Learning.
Classical Hopfield networks [25, 26, 35] have long been recognized for their ability to emulate human brain associative memory, primarily focusing on the storage and retrieval of memory patterns. Originally limited by linear memory capacity [35], these networks now achieve exponential storage capabilities [14] and kernelized [70], allowing for efficient handling of vast memory patterns. Architectural innovations [22, 63, 16] have integrated Hopfield networks into models like Transformers, serving as advanced attention mechanisms [69]. Their biological plausibility [37, 36] has been a key focus, aligning computational models more closely with human associative memory. These developments have expanded applications across various fields, including tabular data learning [73], drug discovery [64], immunology [72], time series forecasting [70], reinforcement learning [58], and large foundation models [31, 16, 20], demonstrating their versatility in addressing complex problems.
Limitations of Transformer Architectures.
While Transformers excel in natural language processing, their capabilities in mathematical computation remain constrained [6]. To understand these limitations, researchers have examined two types of Transformers: (1) average-head attention, which assigns 1 to the largest probability entry and 0 to others, and (2) softmax-attention, which computes probabilities as $\mathsf{Softmax}(X)=\operatorname{diag}(\exp(X)\cdot{\bf 1}_{n})^{-1}\cdot\exp(X)$. Average-head Transformers can recognize languages beyond $\mathsf{AC}^{0}$, but they are confined to $\mathsf{TC}^{0}$, as shown by [57]. Similarly, softmax-attention Transformers are also within $\mathsf{TC}^{0}$ [40]. Extending this, [56] formalize a generalized similarity function $s$ and demonstrate that softmax-attention Transformers fall under $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. Through a first-order logic framework with $\mathsf{MAJORITY}$ quantifiers ($\mathsf{FOM}$[32], [55] show that these Transformers can be simulated in $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$. [7] further enhances precision, proving that Transformers with minimal error ($2^{-O(\mathrm{poly}(n))}$) also belong to this class. For practical problems, [17] reveal that unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$, log-precision Transformers cannot solve arithmetic, equation-solving, or Context-Free Grammar (CFG) membership tasks [61]. These findings explain why Transformers often struggle with mathematical reasoning. Recently, [8] studied the circuit complexity bounds of RoPE-based transformer architectures. Relevant works related to this topic includes looped transformers [3, 49, 9, 38, 50, 10], accelerating attention computation [8, 31, 45, 51, 73, 21, 67, 70, 29, 23, 47, 39, 71, 54, 43, 44, 62, 46, 42, 30], and more related works [74, 66, 15, 65, 18, 52, 41, 28].
Appendix B Background
B.1 Background on Circuit Complexity
Definition B.1 (Boolean circuit).
Definition B.2 (Circuit family and language recognition, Definition 1.10 on page 10 in [68]).
A language $L\subseteq\{0,1\}^{*}$ is a set of binary strings that represent the ”yes” instances of a given decision problem, where $\{0,1\}^{*}$ denotes the set of all finite binary strings of any length.
B.2 Floating-Point Numbers
Definition B.3 (Floating-point number, Definition 9 in [7]).
Definition B.4 (Rounding, Definition 9 in [7]).
Lemma B.5 (Basic floating-point operations in $\mathsf{TC}^{0}$, Lemmas 10, 11 in [7]).
  • Part 1: Let $x_{1},x_{2}$ be two $p$-bits floating point numbers. The arithmetic operations of addition, multiplication, division, and comparison of $x_{1}$ and $x_{2}$, as described in [7], can be executed using a $\mathsf{DLOGTIME}$-uniform threshold circuit characterized by constant depth, and a size bounded by $\operatorname{poly}(n)$. The maximum depth of the circuit required is denoted by $d_{\mathrm{std}}$.
  • Part 2: Let $x_{1},\ldots,x_{n}$ be $n$ p-bit floating-point number. Iterative multiplication of $x_{1},\ldots,x_{n}$ can be executed with a $\mathsf{DLOGTIME}$-uniform threshold circuit characterized by constant depth and a size bounded by $\operatorname{poly}(n)$. The circuit depth for this operation is indicated as $d_{\otimes}$.
  • Part 3: Let $x_{1},\ldots,x_{n}$ be $n$ p-bit floating-point number. Iterative addition of $x_{1},\ldots,x_{n}$, where rounding is applied after the summation is completed, can be achieved using a $\mathsf{DLOGTIME}$-uniform threshold circuit characterized by constant depth, and a size bounded by $\operatorname{poly}(n)$. The depth required for this operation is represented by $d_{\oplus}$.
Lemma B.6 (Computing $\exp$ in $\mathsf{TC}^{0}$, Lemma 12 in [7]).
Lemma B.7 (Approximating square root in $\mathsf{TC}^{0}$, Lemma 12 in [7]).
Appendix C Additional Theoretical Results for Modern Hopfield Model with Chain of Thought
We provide theoretical results for modern Hopfield networks that incorporate the chain of thought (CoT) mechanism. We first describe the architecture of the modern Hopfield model.
C.1 Modern Hopfield Model Architecture with CoT
Let $\mathcal{V}$ be the vocabulary, and consider a Hopfield model with parameters $\theta$ and a maximum input length $n_{\mathrm{max}}$. This model maps an input token sequence $(x_{1},\ldots,x_{n})\in\mathcal{V}^{n}$ (for $n\leq n_{\mathrm{max}}$) to a probability distribution over $\mathcal{V}$, and we denote it by $p_{\Theta}(\cdot|x_{1},\ldots,x_{n})$. We use $\mathsf{MHM}_{\theta}(x)$ as the function that selects the token in $\mathcal{V}$ which can maximize $p_{\Theta}(\cdot|x_{1},\ldots,x_{n})$, defined as:
$\mathsf{MHM}_{\theta}(x_{1},\ldots,x_{n})=\arg\max_{y\in\mathcal{V}}p_{\theta}(y|x_{1},\ldots,x_{n}).$
$\mathsf{MHM}_{\theta}^{i}(x_{1},x_{2},\ldots,x_{n})=\mathsf{MHM}_{\theta}^{i-1}(x_{1},x_{2},\ldots,x_{n},\mathsf{MHM}_{\theta}^{i-1}(x_{1},x_{2},\ldots,x_{n})),$
For every $i\geq 1$, provided $i+n\leq n_{\mathrm{max}}-1$, we have the base case:
$\mathsf{MHM}_{\theta}^{1}(x_{1},x_{2}\ldots,x_{n})=\mathsf{MHM}_{\theta}(x_{1},x_{2},\ldots,x_{n}).$
The token embedding layer parameterized by $\theta_{\mathsf{TE}}\in\mathbb{R}^{d\times|\mathcal{V}|}$ maps tokens in $\mathcal{V}$ to $\mathbb{R}^{d}$: $\theta_{\mathsf{TE}}(x),\quad\forall x\in\mathcal{V}.$ The position encoding layer with parameters $\theta_{\mathsf{PE}}\in\mathbb{R}^{d\times|\mathcal{V}|}$ maps positions in $[n_{\mathrm{max}}]$ to $\mathbb{R}^{d}$: $\theta_{\mathsf{PE}}(n),\quad\forall n\in[n_{\mathrm{max}}].$ The output layer parameterized by $\theta_{\mathsf{OUTPUT}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is defined by:
Algorithm 1 Modern Hopfield Model, $\mathsf{MHM}_{\theta}$ and $p_{\theta}$
3:$h_{i}^{(0)}\leftarrow\theta_{\mathsf{TE}}(x_{i})+\theta_{\mathsf{PE}}(i),\quad\forall i\in[n]$
4:$(h_{1}^{(1)},\ldots,h_{n}^{(1)})\leftarrow\mathsf{MHN}_{\theta_{\mathsf{MHN}}}(h_{1}^{(0)},\ldots,h_{n}^{(0)})$
[…]
6:$\mathsf{MHM}_{\theta}(x)\leftarrow\arg\max_{y}p_{\theta}(y|x_{1},\ldots,x_{n}).$
C.2 Results of Hopfield with CoT
Lemma C.1 (The Hopfield update rule can be interpreted as a form of the transformer attention mechanism, page 5 in [59]).
$Z=\mathrm{softmax}(\frac{1}{\sqrt{d_{k}}}QK^{\top})V=\mathrm{softmax}(\beta\,RW_{Q}W_{K}^{\top}Y^{\top})YW_{K}W_{V},$ (4)
The middle part of Eq. 4 corresponds to the attention mechanism of the transformer. Specifically, when $R=Y$ and $W_{K}W_{V}$ are substituted with $W_{V}$, the Hopfield update rule is reduced to transformer self-attention.
Definition C.2 (Word Problem for Group $G$, [48]).
Lemma C.3 (Theorem 3.5 in [48]).
The proof of Theorem C.4 is a direct consequence of Lemma C.3.
Theorem C.4.
Appendix D Graph Connectivity and Tree Isomorphism Problems
D.1 Undirected Graph Connectivity Problem
Definition D.1 (Undirected graph connectivity problem).
@@ D.1 Undirected Graph Connectivity Problem
Definition D.2 ($\mathsf{FL}$, page 8 on [12]).
@@
Lemma D.4 (Theorem 3 in [11]).
Undirected graph connectivity is $\mathsf{NC}^{1}$-hard for $\mathsf{FL}$ in definition D.2. When the given graph is known to be a disjoint union of cycles, the connectivity problem is $\mathsf{NC}^{1}$-complete for $\mathsf{FL}$.
D.2 Tree Isomorphism Problem
Definition D.5 (Colored trees, page 2 on [34]).
@@ D.2 Tree Isomorphism Problem
Definition D.6 (Isomorphism, page 2 on [34]).
@@
Definition D.7 (Tree isomorphism problem, page 2 on [34]).
@@
Appendix E Proofs of Section 5
E.1 Proofs of Lemma 5.1
Lemma E.1 (Lemma 5.1 Restated).
Assuming that $p\leq\operatorname{poly}(n)$ and the linear affine feature map $\Phi$ is defined as $\Phi(u):=Wu$ for $u,v\in\mathbb{F}_{p}^{d}$, where $W\in\mathbb{F}_{p}^{D_{\Phi}\times d}$ and the linear kernel is $\mathcal{K}(u,v)=u^{\top}W^{\top}Wv$, then the kernelized Hopfield attention matrix $A$ can be simulated using a $\mathsf{DLOGTIME}$-uniform threshold circuit with depth $3d_{\mathrm{std}}+2d_{\oplus}+d_{\mathrm{exp}}$ and size bounded by $\operatorname{poly}(n)$.
@@ Appendix E Proofs of Section 5
We compute the entry of the kernelized attention matrix $A$:
$A_{i,j}:=\exp(\beta\cdot(WR_{i}W_{Q})^{\top}(WY_{j}W_{K})).$
Using Lemma 4.1, the scalar product
$s_{i,j}=(WR_{i}W_{Q})^{\top}(WY_{j}W_{K})$
Next, by Part 1 of Lemma B.5, the product $\beta\cdot s_{i,j}$ requires depth $d_{\mathrm{std}}$.
Finally, by Lemma B.6, the exponential function $\exp(\beta\cdot s_{i,j})$ can be computed with depth $d_{\mathrm{exp}}$ and size $\operatorname{poly}(n)$.
Summing the depths, the overall depth for $A_{i,j}$ is:
$d_{\mathrm{total}}=6d_{\mathrm{std}}+5d_{\oplus}+d_{\mathrm{exp}}.$
E.2 Proofs of Lemma 5.2
Lemma E.2 (Lemma 5.2 Restated).
Assuming that $p\leq\operatorname{poly}(n)$. Then we can simulate kernelized Hopfield layer using a $\mathsf{DLOGTIME}$-uniform threshold circuit with depth $10d_{\mathrm{std}}+8d_{\oplus}+d_{\mathrm{exp}}$ and size bounded by $\operatorname{poly}(n)$.
@@
To compute a single layer, we must multiply the matrices $D^{-1}$, $A$, $Y$, and $W_{V}$. First, we compute $D$ and $A$. Using $D:=\operatorname{diag}(\beta\cdot A1_{n})$, $D$ can be simulated with depth $d_{\oplus}+d_{\mathrm{std}}$ and size $\operatorname{poly}(n)$ by Part 1 and Part 3 of Lemma B.5.
Next, matrix products $A$, $Y$, and $W_{V}$ can be simulated with depth $2(d_{\mathrm{std}}+d_{\oplus})$ and $\operatorname{poly}(n)$ size, as shown in Lemma 4.1. Finally, the computation of $D^{-1}\cdot(AYW_{V})$ can be performed using parallel division, requiring depth $d_{\mathrm{std}}$ and size $\operatorname{poly}(n)$, which follows from Part 1 of Lemma B.5.
Combining these results, the total depth is:
$d_{\mathrm{total}}=10d_{\mathrm{std}}+8d_{\oplus}+d_{\mathrm{exp}}.$
E.3 Proofs of Lemma 5.3
Lemma E.3 (Lemma 5.3 Restated).
Assume that, for each $i\in[m]$, we can simulate the function $f_{i}$ in $\mathsf{KHN}$ by a $\mathsf{DLOGTIME}$-uniform threshold circuit with $\operatorname{poly}(n)$ size and constant depth $d_{f}$. Let the precision $p\leq\operatorname{poly}(n)$. Then we can simulate kernelized Hopfield networks using a $\mathsf{DLOGTIME}$-uniform threshold circuit with depth $(m+1)d_{f}+10md_{\mathrm{std}}+8md_{\oplus}+md_{\mathrm{exp}}$ and size bounded by $\operatorname{poly}(n)$.
@@
$(m+1)d_{g}+10md_{\mathrm{std}}+8md_{\oplus}+md_{\mathrm{exp}}.$
E.4 Proofs of Theorem 5.4
Theorem E.4 (Theorem 5.4 Restated).
Assume that, for each $i\in[m]$, we can simulate the function $f_{i}$ in $\mathsf{KHN}$ by a $\mathsf{DLOGTIME}$-uniform threshold circuit with constant depth $d_{f}$ and size $\operatorname{poly}(n)$. If precision $p\leq\operatorname{poly}(n)$, hidden dimension $d\leq O(n)$, and number of layers $m\leq O(1)$, then we can simulate kernelized Hopfield networks $\mathsf{KHM}$ by a $\mathsf{DLOGTIME}$-uniform circuit family within $\mathsf{TC}^{0}$.
@@
$(m+1)d_{g}+10md_{\mathrm{std}}+8md_{\oplus}+md_{\mathrm{exp}}=O(1),$

### v2

@@ 4.2 Computing Hopfield Attention Matrix
    The term $\beta\cdot s_{i,j}$ is computed with depth $d_{\mathrm{std}}$ using Part 1 of Lemma 3.3.
@@
    Using Lemma 3.4, the exponential function $\exp(\beta\cdot s_{i,j})$ is computable with depth $d_{\mathrm{exp}}$.
Summing the circuit depths for all steps, the total depth required for $A_{i,j}$ is $d_{\mathrm{total}}=4d_{\mathrm{std}}+3d_{\oplus}+d_{\mathrm{exp}}.$ Since all entries of $A$ can be simulated in parallel, the total depth $4d_{\mathrm{std}}+3d_{\oplus}+d_{\mathrm{exp}}$ and size $\operatorname{poly}(n)$, completing the proof. ∎
@@ 4.3 Computing Single Hopfield Layer
The computation involves multiplying matrices $D^{-1},A,Y,\widetilde{W}_{V}$. First, we compute $D^{-1}$ and $A$:
Part 1. $D=\mathrm{diag}(\beta\cdot A\mathbf{1}_{n})$ is computed with depth $d_{\oplus}+d_{\mathrm{std}}$, as shown in Part 1 and Part 3 of Lemma 3.3. If we compute $D^{-1}$ , it requires a depth of $d_{\oplus}+2d_{\mathrm{std}}$ follows from Part 1 of Lemma 3.3.
Part 2. From Lemma 4.4, $A$ is computed with depth $4d_{\mathrm{std}}+3d_{\oplus}+d_{\mathrm{exp}}$.
Part 3. Then, we can multiply $A,Y$ and $\widetilde{W}_{V}$, which can be computed by a depth $2(d_{\mathrm{std}}+d_{\oplus})$, size $\operatorname{poly}(n)$ uniform threshold circuit following from Lemma 4.3.
Part 4. Finally, the division $D^{-1}\cdot(AY\widetilde{W}_{V})$ is computed in parallel with depth $d_{\mathrm{std}}$. Combining these depths, $d_{\mathrm{total}}=8d_{\mathrm{std}}+6d_{\oplus}+d_{\mathrm{exp}}.$ The total size of the circuit remains $\operatorname{poly}(n)$, completing the proof. ∎
@@ 4.4 Computing Common Components Layers
For each $i\in[n]$, Lemma 4.3 guarantees that the computation of $W_{1}\cdot X_{i,*}$ requires a $\mathsf{DLOGTIME}$-uniform circuit which has depth $d_{\mathrm{std}}+d_{\oplus}$ and size $\operatorname{poly}(n)$. Applying Part 1 of Lemma 3.3, we can compute $W_{1}\cdot X_{i,*}+b_{1}$ with an additional $\mathsf{DLOGTIME}$-uniform circuit which has depth $d_{\mathrm{std}}$ and size $\operatorname{poly}(n)$. The ReLU activation, $\mathsf{ReLU}(W_{1}\cdot X_{i,*}+b_{1})$, also requires depth $d_{\mathrm{std}}$ and size $\operatorname{poly}(n)$ as per Part 1 of Lemma 3.3.
@@ 4.6 Main Result: Circuit Complexity Bound of Modern Hopfield Networks
Theorem 4.8 implies that unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$, Modern Hopfield Networks with $\operatorname{poly}(n)$-precision, constant-depth, $\operatorname{poly}(n)$-size can be simulated by a $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ circuit family. It means that although Modern Hopfield Networks gain success empirically, it still suffers fundamental expressivity limitations under circuit complexity.
@@ 5 Complexity of Kernelized Hopfield Networks
In this section, we extend the complexity results of Modern Hopfield Networks to Kernelized Hopfield Networks. The formal definition of Kernelized Hopfield Networks is provided in Section 3.4.
Section 5.1 analyzes the computation of the kernelized Hopfield attention matrix. In Section 5.2, we examine the computation of a single kernelized Hopfield layer. Section 5.3 details the computation of the entire kernelized Hopfield network. Finally, in Section 5.4, we present the main results on the circuit complexity bounds for Kernelized Hopfield Networks.
@@ 5.1 Computing Kernelized Hopfield Attention Matrix
We compute the entry of the kernelized attention matrix $A$:
$A_{i,j}:=\exp(\beta\cdot(WR_{i}W_{Q})^{\top}(WY_{j}W_{K})).$
Using Lemma 4.3, the scalar product
$s_{i,j}=(WR_{i}W_{Q})^{\top}(WY_{j}W_{K})$
Next, by Part 1 of Lemma 3.3, the product $\beta\cdot s_{i,j}$ requires depth $d_{\mathrm{std}}$.
Finally, by Lemma 3.4, the exponential function $\exp(\beta\cdot s_{i,j})$ can be computed with depth $d_{\mathrm{exp}}$ and size $\operatorname{poly}(n)$.
Summing the depths, the overall depth for $A_{i,j}$ is:
$d_{\mathrm{total}}=6d_{\mathrm{std}}+5d_{\oplus}+d_{\mathrm{exp}}.$
@@ 5.2 Computing Single Kernelized Hopfield Layer
To compute a single layer, we must multiply the matrices $D^{-1}$, $A$, $Y$, and $W_{V}$. First, we compute $D$ and $A$. Using $\operatorname{diag}(\beta\cdot A1_{n}), DD$ can be simulated with depth $d_{\oplus}+d_{\mathrm{std}}$ and size $\operatorname{poly}(n)$ by Part 1 and Part 3 of Lemma 3.3.
Next, matrix products $A$, $Y$, and $W_{V}$ can be simulated with depth $2(d_{\mathrm{std}}+d_{\oplus})$ and $\operatorname{poly}(n)$ size, as shown in Lemma 4.3. Finally, the computation of $D^{-1}\cdot(AYW_{V})$ can be performed using parallel division, requiring depth $d_{\mathrm{std}}$ and size $\operatorname{poly}(n)$, which follows from Part 1 of Lemma 3.3.
Combining these results, the total depth is:
$d_{\mathrm{total}}=10d_{\mathrm{std}}+8d_{\oplus}+d_{\mathrm{exp}}.$
@@ 5.3 Computing Kernelized Hopfield Networks
$(m+1)d_{g}+10md_{\mathrm{std}}+8md_{\oplus}+md_{\mathrm{exp}}.$
@@ 5.4 Main Result: Circuit Complexity Bound of Kernelized Hopfield Networks
$(m+1)d_{g}+10md_{\mathrm{std}}+8md_{\oplus}+md_{\mathrm{exp}}=O(1),$
@@ 6 Hardness
In this section, we present two computational problems along with their corresponding hardness results. In Section 6.1, we introduce the undirected graph connectivity problem. In Section 6.2, we introduce the tree isomorphism problem. In Section 6.3, we outline the corresponding hardness results.
6.1 Undirected Graph Connectivity Problem
Definition 6.1 (Undirected graph connectivity problem).
@@ 6.1 Undirected Graph Connectivity Problem
Definition 6.2 ($\mathsf{FL}$, page 8 on [9]).
@@
Lemma 6.4 (Theorem 3 in [8]).
Undirected graph connectivity is $\mathsf{NC}^{1}$-hard for $\mathsf{FL}$ in definition 6.2. When the given graph is known to be a disjoint union of cycles, the connectivity problem is $\mathsf{NC}^{1}$-complete for $\mathsf{FL}$.
6.2 Tree Isomorphism Problem
Definition 6.5 (Colored trees, page 2 on [27]).
@@ 6.2 Tree Isomorphism Problem
Definition 6.6 (Isomorphism, page 2 on [27]).
@@
Definition 6.7 (Tree isomorphism problem, page 2 on [27]).
@@
6.3 Hardness Results
Theorem 6.9.
@@ 6.3 Hardness Results
This result is obtained by integrating Theorem 4.8 (the circuit complexity bound of Modern Hopfield Networks), Lemma 6.4 (showing that undirected graph connectivity is $\mathsf{NC}^{1}$-complete), and Fact 6.3. ∎
Theorem 6.10.
@@
This result derives from Theorem 4.8 and Lemma 6.8 (showing that tree isomorphism is $\mathsf{NC}^{1}$-complete). ∎
Theorem 6.11.
@@
This result is a direct consequence of Theorem 5.4 and Lemma 6.4. ∎
Theorem 6.12.
@@
This result is a direct consequence of Theorem 5.4 and Lemma 6.8. ∎
7 Modern Hopfield Model with Chain of Thought
We provide theoretical results for Modern Hopfield Networks that incorporate the Chain-of-Thought (CoT) mechanism. We first describe the architecture of the modern Hopfield model.
7.1 Modern Hopfield Model Architecture with CoT
Let $\mathcal{V}$ be the vocabulary, and consider a Hopfield model with parameters $\theta$ and a maximum input length $n_{\mathrm{max}}$. This model maps an input token sequence $(x_{1},\ldots,x_{n})\in\mathcal{V}^{n}$ (for $n\leq n_{\mathrm{max}}$) to a probability distribution over $\mathcal{V}$, and we denote it by $p_{\Theta}(\cdot|x_{1},\ldots,x_{n})$. We use $\mathsf{MHM}_{\theta}(x)$ as the function that selects the token in $\mathcal{V}$ which can maximize $p_{\Theta}(\cdot|x_{1},\ldots,x_{n})$, defined as: $\mathsf{MHM}_{\theta}(x_{1},\ldots,x_{n}):=\arg\max_{y\in\mathcal{V}}p_{\theta}(y|x_{1},\ldots,x_{n}).$
$\mathsf{MHM}_{\theta}^{i}(x_{1},x_{2},\ldots,x_{n})=\mathsf{MHM}_{\theta}^{i-1}(x_{1},x_{2},\ldots,x_{n},\mathsf{MHM}_{\theta}^{i-1}(x_{1},x_{2},\ldots,x_{n})).$
For every $i\geq 1$, provided $i+n\leq n_{\mathrm{max}}-1$, we have the base case: $\mathsf{MHM}_{\theta}^{1}(x_{1},x_{2}\ldots,x_{n})=\mathsf{MHM}_{\theta}(x_{1},x_{2},\ldots,x_{n}).$
The token embedding layer parameterized by $\theta_{\mathsf{TE}}\in\mathbb{R}^{d\times|\mathcal{V}|}$ maps tokens in $\mathcal{V}$ to $\mathbb{R}^{d}$: $\theta_{\mathsf{TE}}(x),\quad\forall x\in\mathcal{V}.$ The position encoding layer with parameters $\theta_{\mathsf{PE}}\in\mathbb{R}^{d\times|\mathcal{V}|}$ maps positions in $n_{\mathrm{max}}]$ to $\mathbb{R}^{d}$: $\theta_{\mathsf{PE}}(n),\quad\forall n\in[n_{\mathrm{max}}].$ The output layer parameterized by $\theta_{\mathsf{OUTPUT}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is defined by:
Algorithm 1 Modern Hopfield Model, $\mathsf{MHM}_{\theta}$ and $p_{\theta}$
[…]
7.2 Results of Hopfield with CoT
Lemma 7.1 (The Hopfield update rule can be interpreted as a form of the transformer attention mechanism, page 5 in [40]).
$Z=$ $~\mathrm{softmax}(\frac{1}{\sqrt{d_{k}}}QK^{\top})V$
$=$ $~\mathrm{softmax}(\beta\,RW_{Q}W_{K}^{\top}Y^{\top})YW_{K}W_{V},$ (4)
The middle part of Eq. 7.1 corresponds to the attention mechanism of the transformer. Specifically, when $R=Y$ and $W_{K}W_{V}$ are substituted with $W_{V}$, the Hopfield update rule is reduced to transformer self-attention.
Definition 7.2 (Word Problem for Group $G$, [33]).
Lemma 7.3 (Theorem 3.5 in [33]).
The proof of Theorem 7.4 is a direct consequence of Lemma 7.3.
Theorem 7.4.
8 Conclusion
In this study, we conduct a comprehensive theoretical investigation of Modern Hopfield Networks and Kernelized Hopfield Networks, establishing key limitations on their computational power. Our analysis focuses on the circuit complexity of individual architectural components, showing that these networks can be simulated using uniform $\mathsf{TC}^{0}$ circuits. Crucially, we prove that unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$, both modern and Kernelized Hopfield Networks, when implemented with polynomial precision, a constant number of layers, and hidden dimensions satisfying $d\leq O(n)$, are unable to solve problems such as undirected graph connectivity and tree isomorphism. However, we demonstrate that this theoretical ceiling is not insurmountable: by incorporating a Chain-of-Thought mechanism, Modern Hopfield Networks can transcend the $\mathsf{TC}^{0}$ barrier and successfully solve such $\mathsf{NC}^{1}$-hard problems.